OpenStudy (anonymous):

What is the measure of the angle formed by the minute hand and the hour hand of a clock when the clock shows 4:35? A. 72.5° B. 90° C. 120° D. 72°

OpenStudy (anonymous):

Lets take 4 o clock as the reference point for both the arms of he clock

OpenStudy (anonymous):

One way is to find the angle of each hand in relation to 12 o'clock (0 degrees)

OpenStudy (anonymous):

Convert the min portion to decimal hrs: 4.35 = 4.833 Find the position of the hr hand: 4.833 * 30 = 137.5 degrees Find the position of the min hand: 35 * 6 = 210 degrees Find the difference: 210 - 137.5 = 72.5 degree angle formed at 3:35 c. 11:20 Decimal hrs: 11:20 = 11.333 hr hand: 11.333 * 30 = 340 degrees Min hand: 20 * 6 =120 degrees